OpenStudy (anonymous):

if the radius of circle G is 10, and the measure of angle AGB if 72 degrees , calculate the area of the sector in circle G formed by the segments AG and GB

OpenStudy (campbell_st):

won't it simply be \[Area = \frac{x}{360} \times \pi \times r^2\] where x is the sector angle... I think your sector angle is 72

OpenStudy (anonymous):

yeah! thanks!

OpenStudy (campbell_st):

Oh, and AG and BG are both radii...